# On the Design of Expressive and Trainable Pulse-based Quantum Machine Learning Models

### v1

@@ A.1 Model 1
where $\mathfrak{g}=\{iJ_{\rm x},iJ_{\rm y},iJ_{\rm z}\}_{LA}={\mathfrak{su}}(2)$ is the $2^{n}$-dimensional unitary irreducible representation of ${\mathfrak{su}}(2)$. Let $J_{\pm}=J_{\rm x}\pm iJ_{\rm y}$ be the raising and lowering operators, and
@@
be the basis states. Under the basis state, we have
@@
Since the maximum and minimum eigenvalues of $J_{\rm z}$ are $\pm\frac{2^{n}-1}{2}$, and the range of the target function is chosen to be [1,1][-1,1], we normalize the observable to $M=\frac{2}{2^{n}-1}J_{\rm z}$. According to Theorem 1, it is easy to obtain that
@@
If the initial state $|\psi_{0}\rangle=|0\rangle^{\otimes n}$, it can be verified that the model satisfies the condition for expressivity.
On the other hand, the model is uncontrollable and $\rm dim(\mathfrak{g})=3$. Using the orthonormal basis
@@
of the Lie algebra ${\mathfrak{su}}(2)$, we have
${\rm Var}\left[f(\textbf{x},\Theta)\right]$
$=$ $\frac{1}{3}P_{\mathfrak{g}}(\rho)P_{\mathfrak{g}}(M)$
@@ A.2 Model 2
The Lie algebra generated by this model is $\mathfrak{g}=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{n}$, where each $\mathfrak{g}_{j}={\mathfrak{su}}(2)$, thus $\dim(\mathfrak{g})=3n$. In this case,
$\mathcal{S}_{k}=\{\sigma_{\rm x}^{(i)},\sigma_{\rm y}^{(i)},\sigma_{\rm z}^{(i)}\}_{1\leq i\leq n},k\in N.$
@@
Given that the orthonormal basis of $\mathfrak{g}_{j}$ is $i\{2^{-n/2}\sigma_{\rm x}^{(j)},2^{-n/2}\sigma_{\rm y}^{(j)},2^{-n/2}\sigma_{\rm z}^{(j)}\}$, we find $M_{\mathfrak{g}_{j}}=\frac{1}{n}\sigma_{\rm z}^{(j)}$, and $\rho_{\mathfrak{g}_{j}}=2^{-n}\sigma_{\rm z}^{(j)}$ for all $j$. Then we obtain
@@ A.3 Model 3
We can derive that
$\mathcal{S}_{k}$ $=$ $\{\widehat{\sigma_{\rm x}^{(i)}\sigma_{\rm y}^{(j)}}\}_{1\leq i<j\leq n},k\in N,$
where $\widehat{\sigma_{\rm x}^{(i)}\sigma_{\rm y}^{(j)}}=\sigma_{\rm x}^{(i)}\sigma_{\rm z}^{(i+1)}\cdots\sigma_{\rm z}^{(j-1)}\sigma_{\rm y}^{(j)}$. For the selected initial state $|\psi_{0}\rangle=|0\rangle^{\otimes(n-2)}\otimes(\frac{1}{\sqrt{2}}|0\rangle+\frac{1}{\sqrt{2}}|1\rangle)\otimes(\frac{1}{\sqrt{2}}|0\rangle+\frac{i}{\sqrt{2}}|1\rangle)$, we can verify that for any $n\geq 3$, this model satisfies the condition in Theorem 1.
As for the trainability of the model, the Lie algebra associated to this model is $\mathfrak{g}={\mathfrak{so}}(n)$ that has dimension $\dim(\mathfrak{g})=\frac{n(n-1)}{2}\in\mathcal{O}(\text{poly}(n))$. Since the orthonormal basis of ${\mathfrak{so}}(n)$ is
@@
we can easily get $P_{\mathfrak{g}}(M)=2^{n}$ and $P_{\mathfrak{g}}(\rho)=2^{-n}$. Then we have
@@ A.4 Model 4
The model can be verified to possess full expressivity according to the criterion proposed in [22] because the control Hamiltonians generate the entire $\mathfrak{su}(2^{n})$ Lie algebra.
For the calculation of variance, based on the analysis of [41], we have

### v2

@@ A.1 Model 1
where $\mathfrak{g}=\{iJ_{\rm x},iJ_{\rm y},iJ_{\rm z}\}_{LA}={\mathfrak{su}}(2)$ is the $2^{n}$-dimensional unitary irreducible representation of ${\mathfrak{su}}(2)$. Let $J_{\pm}=J_{\rm x}\pm iJ_{\rm y}$ denote the raising and lowering operators, and
@@
the basis states. In this basis:
@@
Since the maximum and minimum eigenvalues of $J_{\rm z}$ are $\pm\frac{2^{n}-1}{2}$, and the target function range is [1,1][-1,1], we normalize the observable as $M=\frac{2}{2^{n}-1}J_{\rm z}$. From Theorem 1:
@@
With initial state $|\psi_{0}\rangle=|0\rangle^{\otimes n}$, the model satisfies the expressivity condition.
The model is uncontrollable with $\rm dim(\mathfrak{g})=3$. Using the orthonormal basis
@@
for $\mathfrak{su}(2)$, we obtain:
${\rm Var}\left[f(\textbf{x},\Theta)\right]$ $=$ $\frac{1}{3}P_{\mathfrak{g}}(\rho)P_{\mathfrak{g}}(M)$
@@ A.2 Model 2
The generated Lie algebra is $\mathfrak{g}=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{n}$, where each $\mathfrak{g}_{j}={\mathfrak{su}}(2)$, giving $\dim(\mathfrak{g})=3n$. In this case:
$\mathcal{S}_{k}=\{\sigma_{\rm x}^{(i)},\sigma_{\rm y}^{(i)},\sigma_{\rm z}^{(i)}\}_{1\leq i\leq n}\text{ for }k\in N.$
@@
The orthonormal basis for $\mathfrak{g}_{j}$ is $i\{2^{-n/2}\sigma_{\rm x}^{(j)},2^{-n/2}\sigma_{\rm y}^{(j)},2^{-n/2}\sigma_{\rm z}^{(j)}\}$, and we have $M_{\mathfrak{g}_{j}}=\frac{1}{n}\sigma_{\rm z}^{(j)}$ and $\rho_{\mathfrak{g}_{j}}=2^{-n}\sigma_{\rm z}^{(j)}$ for all $j$. Then:
@@ A.3 Model 3
We derive:
$\mathcal{S}_{k}$ $=$ $\{\widehat{\sigma_{\rm x}^{(i)}\sigma_{\rm y}^{(j)}}\}_{1\leq i<j\leq n}\text{ for }k\in N,$
where $\widehat{\sigma_{\rm x}^{(i)}\sigma_{\rm y}^{(j)}}=\sigma_{\rm x}^{(i)}\sigma_{\rm z}^{(i+1)}\cdots\sigma_{\rm z}^{(j-1)}\sigma_{\rm y}^{(j)}$. With initial state $|\psi_{0}\rangle=|0\rangle^{\otimes(n-2)}\otimes(\frac{1}{\sqrt{2}}|0\rangle+\frac{1}{\sqrt{2}}|1\rangle)\otimes(\frac{1}{\sqrt{2}}|0\rangle+\frac{i}{\sqrt{2}}|1\rangle)$, we verify that for $n\geq 3$, the model satisfies Theorem 1.
The Lie algebra is $\mathfrak{g}={\mathfrak{so}}(n)$ with dimension $\dim(\mathfrak{g})=\frac{n(n-1)}{2}\in\mathcal{O}(\text{poly}(n))$. The orthonormal basis for ${\mathfrak{so}}(n)$ is
@@
giving $P_{\mathfrak{g}}(M)=2^{n}$ and $P_{\mathfrak{g}}(\rho)=2^{-n}$. Then:
@@ A.4 Model 4
The model possesses full expressivity according to the criterion in [22] since the control Hamiltonians generate the full $\mathfrak{su}(2^{n})$ Lie algebra.
For variance calculation, based on [41]: